# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
Appendix I Proof of Proposition 5.1
@@
For any slice $u_{i}\in\mathcal{U}$, this implies
$\bigl|\langle\bar{g}_{t}(\mathcal{B}(x;r))-\bar{g}_{t}(\mathcal{B}(x^{\prime};r)),u_{i}\rangle\bigr|\leq\frac{2C}{B_{t}}K_{t}(x,x^{\prime};r).$ (34)
@@
By Assumption 5.2, for each $u_{i}\in\mathcal{U}$ there exists $L_{t,i}<\infty$ such that, for all $z,z^{\prime}\in\mathbb{R}^{d}$ and all histories $y_{<t}$,
@@
$\bigl|\langle f_{t}(x,y_{<t};r)-f_{t}(x^{\prime},y_{<t};r),u_{i}\rangle\bigr|$ $\leq L_{t,i}\,\frac{2C}{B_{t}}\,K_{t}(x,x^{\prime};r)$
@@
$\bigl|\langle f_{t}(x,y_{<t};r)-f_{t}(x^{\prime},y_{<t};r),u_{i}\rangle\bigr|^{2}\leq\Bigl(\frac{2CL_{t,i}}{B_{t}}K_{t}(x,x^{\prime};r)\Bigr)^{2}.$
@@
$\bigl|\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{i}\rangle\bigr|^{2}\leq\Bigl(\frac{2K_{t}L_{t,i}C}{B_{t}}\Bigr)^{2}=:h_{t,i}.$
@@
Appendix J Proof of Proposition 5.2
Fix an iteration $t$ and a slice profile $\{\mathcal{U},\omega\}$ with $\mathcal{U}=\{u_{i}\}_{i=1}^{m}\subset\mathbb{S}^{d-1}$. Let $\theta\in\Theta$, $(s_{i},s_{j})\in\mathcal{Q}$, a history $y_{<t}$ in the support, and a coupling $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$ be arbitrary. Let $(X,X^{\prime})\sim\gamma$. In addition, let $\eta_{t}\sim\mathbb{P}_{\eta}$ denote the subsampling randomness at iteration $t$. We write $\bar{g}_{t}(X;\eta_{t})$ and $\bar{g}_{t}(X^{\prime};\eta_{t})$ for the clipped, averaged minibatch gradients, and $f_{t}(X,y_{<t};\eta_{t})$, $f_{t}(X^{\prime},y_{<t};\eta_{t})$ for the corresponding pre-perturbation updates.
Step 1. By per-example $\ell_{2}$-clipping at threshold $C$ and the definition of the discrepancy $K_{t}(X,X^{\prime};\eta_{t})$, the argument in the proof of Proposition 5.1 gives, for every realization of $\eta_{t}$,
$\bigl\|\bar{g}_{t}(X;\eta_{t})-\bar{g}_{t}(X^{\prime};\eta_{t})\bigr\|_{2}\leq\frac{2C}{B_{t}}K_{t}(X,X^{\prime};\eta_{t}).$
Hence, for any slice $u_{i}\in\mathcal{U}$,
$\bigl|\langle\bar{g}_{t}(X;\eta_{t})-\bar{g}_{t}(X^{\prime};\eta_{t}),u_{i}\rangle\bigr|\leq\frac{2C}{B_{t}}K_{t}(X,X^{\prime};\eta_{t}).$ (35)
@@
$\Delta_{t}^{\mathrm{pre}}(X,X^{\prime};\eta_{t}):=\bar{g}_{t}(X;\eta_{t})-\bar{g}_{t}(X^{\prime};\eta_{t}),$
@@
$\Delta_{t}(X,X^{\prime};\eta_{t}):=f_{t}(X,y_{<t};\eta_{t})-f_{t}(X^{\prime},y_{<t};\eta_{t}).$
By Assumption 5.2, for each slice $u_{i}\in\mathcal{U}$ and every realization of $\eta_{t}$,
$\bigl|\langle\Delta_{t}(X,X^{\prime};\eta_{t}),u_{i}\rangle\bigr|\leq L_{t,i}\bigl|\langle\Delta_{t}^{\mathrm{pre}}(X,X^{\prime};\eta_{t}),u_{i}\rangle\bigr|.$
@@
$\bigl|\langle\Delta_{t}(X,X^{\prime};\eta_{t}),u_{i}\rangle\bigr|\leq\frac{2L_{t,i}C}{B_{t}}\,K_{t}(X,X^{\prime};\eta_{t})$ (36)
@@
Step 2. Fix $(X,X^{\prime})$ in the support. Squaring both sides of (36) and taking expectations with respect to $\eta_{t}\sim\mathbb{P}_{\eta}$, we obtain, for each $i=1,\dots,m$,
$\mathbb{E}_{\eta_{t}}\bigl[\bigl|\langle\Delta_{t}(X,X^{\prime};\eta_{t}),u_{i}\rangle\bigr|^{2}\bigr]\leq\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\mathbb{E}_{\eta_{t}}\bigl[K_{t}(X,X^{\prime};\eta_{t})^{2}\bigr].$
@@
$\mathbb{E}_{\eta_{t}}\bigl[K_{t}(X,X^{\prime};\eta_{t})^{2}\bigr]\leq\overline{K}_{t}^{2}.$
@@
$\mathbb{E}_{\eta_{t}}\bigl[\bigl|\langle\Delta_{t}(X,X^{\prime};\eta_{t}),u_{i}\rangle\bigr|^{2}\bigr]\leq\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2}=h_{t,i}^{\mathsf{ms}},\quad i=1,\dots,m.$
@@
$h_{t}^{\mathsf{ms}}=\bigl(h_{t,i}^{\mathsf{ms}}\bigr)_{i=1}^{m}\quad\textup{and}\quad h_{t,i}^{\mathsf{ms}}=\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2},$
is a valid ms-HUC at iteration $t$. ∎
Appendix K Proof of Lemma 5.1

### v3

@@ Appendix H Minimality and Attainability of HUC
Appendix N Proof of Theorem 4.7
Let $U_{1:m}\overset{\mathrm{iid}}{\sim}\omega$. Define the random variables
$X_{\ell}:=\big(\Delta_{\infty}^{U_{\ell}}\big)^{2}\in[0,\Delta_{0}^{2}],\qquad\ell=1,\ldots,m,$
and recall
$\overline{\Delta}^{2}(\omega):=\mathbb{E}_{U\sim\omega}\!\big[(\Delta_{\infty}^{U})^{2}\big]=\mathbb{E}[X_{1}].$
Step 1.
Define the event
$\mathcal{E}_{\mathrm{dist}}:=\Big\{\Delta_{\infty}^{U_{\ell}}\leq\widehat{\Delta}_{\infty}(U_{\ell})\ \text{for all }\ell\in[m]\Big\}.$
By assumption,
$\Pr(\mathcal{E}_{\mathrm{dist}})\geq 1-\gamma/2.$ (57)
On $\mathcal{E}_{\mathrm{dist}}$,
$\frac{1}{m}\sum_{\ell=1}^{m}X_{\ell}=\frac{1}{m}\sum_{\ell=1}^{m}\big(\Delta_{\infty}^{U_{\ell}}\big)^{2}\leq\frac{1}{m}\sum_{\ell=1}^{m}\big(\widehat{\Delta}_{\infty}(U_{\ell})\big)^{2}.$ (58)
Step 2.
Since $X_{1},\ldots,X_{m}$ are i.i.d. and bounded in $0,\Delta_{0}^{2}]$, Hoeffding’s inequality yields that for any $\eta\in(0,1)$,
$\Pr\!\left(\mathbb{E}[X_{1}]\leq\frac{1}{m}\sum_{\ell=1}^{m}X_{\ell}+\Delta_{0}^{2}\sqrt{\frac{\log(2/\eta)}{2m}}\right)\geq 1-\eta.$
Setting $\eta=\gamma/2$ and defining the event
$\mathcal{E}_{\mathrm{mc}}:=\left\{\overline{\Delta}^{2}(\omega)\leq\frac{1}{m}\sum_{\ell=1}^{m}X_{\ell}+\Delta_{0}^{2}\sqrt{\frac{\log(4/\gamma)}{2m}}\right\},$
we obtain
$\Pr(\mathcal{E}_{\mathrm{mc}})\geq 1-\gamma/2.$ (59)
Step 3.
On the intersection event $\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}}$, combining the defining inequality of $\mathcal{E}_{\mathrm{mc}}$ with (58) gives
$\overline{\Delta}^{2}(\omega)\leq\frac{1}{m}\sum_{\ell=1}^{m}\big(\widehat{\Delta}_{\infty}(U_{\ell})\big)^{2}+\Delta_{0}^{2}\sqrt{\frac{\log(4/\gamma)}{2m}}.$ (60)
By the union bound together with (57) and (59),
$\Pr(\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}})\geq 1-\gamma.$ (61)
Step 4.
Assume the mechanism uses $N\sim\mathcal{N}(0,\sigma^{2}I_{d})$ with
$\sigma^{2}\;\geq\;\frac{\alpha}{2\varepsilon}\left(\frac{1}{m}\sum_{\ell=1}^{m}\big(\widehat{\Delta}_{\infty}(U_{\ell})\big)^{2}+\Delta_{0}^{2}\sqrt{\frac{\log(4/\gamma)}{2m}}\right).$
On $\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}}$, inequality (60) implies
$\sigma^{2}\;\geq\;\frac{\alpha}{2\varepsilon}\,\overline{\Delta}^{2}(\omega).$
Therefore, by Theorem 4.5 (Gaussian calibration for $(\alpha,\varepsilon,\omega)$-Ave-SRPP), the mechanism $\mathcal{M}(X)=f(X)+N$ satisfies $(\alpha,\varepsilon,\omega)$-Ave-SRPP on this event. Finally, (61) gives that this conclusion holds with probability at least $1-\gamma$ over both the conditional sampling (used to form $\widehat{\Delta}_{\infty}$) and the draw of $U_{1:m}$. ∎
Appendix O Proof of Theorem 4.8
Fix $\sigma^{2}>0$ and abbreviate $c:=c(\sigma)=\frac{\alpha(\alpha-1)}{2\sigma^{2}}$ and $b:=b(\sigma)=\exp(c\Delta_{0}^{2})$. Let $U_{1:m}\overset{\mathrm{iid}}{\sim}\omega$ and define the random variables
$Y_{\ell}\;:=\;\exp\!\big(c\,(\Delta_{\infty}^{U_{\ell}})^{2}\big),\qquad\ell=1,\ldots,m.$
Since $0\leq\Delta_{\infty}^{u}\leq\Delta_{0}$ for all $u$, we have $Y_{\ell}\in[1,b]$ almost surely, and
$\mu(\sigma)\;:=\;\mathbb{E}_{U\sim\omega}\!\Big[\exp\!\big(c\,(\Delta_{\infty}^{U})^{2}\big)\Big]\;=\;\mathbb{E}[Y_{1}].$
Step 1.
Define the event
$\mathcal{E}_{\mathrm{dist}}:=\Big\{\Delta_{\infty}^{U_{\ell}}\leq\widehat{\Delta}_{\infty}(U_{\ell})\ \text{for all }\ell\in[m]\Big\}.$
By assumption, $\Pr(\mathcal{E}_{\mathrm{dist}})\geq 1-\gamma/2$. On $\mathcal{E}_{\mathrm{dist}}$, for each $\ell$,
$Y_{\ell}=\exp\!\big(c\,(\Delta_{\infty}^{U_{\ell}})^{2}\big)\;\leq\;\exp\!\big(c\,(\widehat{\Delta}_{\infty}(U_{\ell}))^{2}\big),$
and hence
$\frac{1}{m}\sum_{\ell=1}^{m}Y_{\ell}\;\leq\;\widehat{\mu}_{m}(\sigma):=\frac{1}{m}\sum_{\ell=1}^{m}\exp\!\Big(c\,\big(\widehat{\Delta}_{\infty}(U_{\ell})\big)^{2}\Big).$ (62)
Step 2.
Since $Y_{1},\ldots,Y_{m}$ are i.i.d. and bounded in $[1,b]$, by Hoeffding’s inequality, we have, for any $\eta\in(0,1)$,
$\Pr\!\left(\mu(\sigma)\leq\frac{1}{m}\sum_{\ell=1}^{m}Y_{\ell}+(b-1)\sqrt{\frac{\log(2/\eta)}{2m}}\right)\geq 1-\eta.$
Setting $\eta=\gamma/2$ gives the stated $\log(4/\gamma)$ term. Set $\eta=\gamma/2$ and define the event
$\mathcal{E}_{\mathrm{mc}}:=\left\{\mu(\sigma)\leq\frac{1}{m}\sum_{\ell=1}^{m}Y_{\ell}+(b-1)\sqrt{\frac{\log(4/\gamma)}{2m}}\right\}.$
Then $\Pr(\mathcal{E}_{\mathrm{mc}})\geq 1-\gamma/2$.
Step 3.
On $\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}}$, combining the defining inequality of $\mathcal{E}_{\mathrm{mc}}$ with (62) yields
$\mu(\sigma)\leq\widehat{\mu}_{m}(\sigma)+(b-1)\sqrt{\frac{\log(4/\gamma)}{2m}}.$ (63)
Moreover, by the union bound,
$\Pr(\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}})\;\geq\;1-\gamma.$
Step 4.
For Gaussian SWM with noise $N\sim\mathcal{N}(0,\sigma^{2}I_{d})$, the joint envelope equals
$\mathtt{JR}^{\infty}_{\alpha,\omega}=\frac{1}{\alpha-1}\log\mu(\sigma).$
On $\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}}$, inequality (63) implies
$\mathtt{JR}^{\infty}_{\alpha,\omega}\leq\frac{1}{\alpha-1}\log\!\left(\widehat{\mu}_{m}(\sigma)+(b-1)\sqrt{\frac{\log(4/\gamma)}{2m}}\right).$
Hence, if $\sigma^{2}$ satisfies the stated condition
$\frac{1}{\alpha-1}\log\!\left(\widehat{\mu}_{m}(\sigma)+(b-1)\sqrt{\frac{\log(4/\gamma)}{2m}}\right)\leq\varepsilon,$
then on $\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}}$ we have $\mathtt{JR}^{\infty}_{\alpha,\omega}\leq\varepsilon$. Therefore, by the Gaussian Joint-SRPP calibration result (Theorem 4.6 stated for SRPP), the mechanism $\mathcal{M}(X)=f(X)+\mathcal{N}(0,\sigma^{2}I_{d})$ satisfies $(\alpha,\varepsilon,\omega)$-Joint-SRPP on this event. Since $\Pr(\mathcal{E}_{\mathrm{dist}}\cap\mathcal{E}_{\mathrm{mc}})\geq 1-\gamma$, the claim follows. ∎
Appendix P Formal Version and Proof of Proposition 4.1
We compare the cost of evaluating (i) sliced sensitivities $\Delta_{\infty}^{u}$ over finitely many directions versus (ii) the unsliced $d$-dimensional empirical $W_{\infty}$ sensitivity.
[…]
Lemma P.1 (Exact empirical $W_{\infty}$ on $\mathbb{R}$).
Let $a_{1},\ldots,a_{n}\in\mathbb{R}$ and $b_{1},\ldots,b_{n}\in\mathbb{R}$ and define $\widehat{P}:=\frac{1}{n}\sum_{k=1}^{n}\delta_{a_{k}}$ and $\widehat{Q}:=\frac{1}{n}\sum_{k=1}^{n}\delta_{b_{k}}$. Let $a_{(1)}\leq\cdots\leq a_{(n)}$ and $b_{(1)}\leq\cdots\leq b_{(n)}$ be the order statistics. Then
$W_{\infty}(\widehat{P},\widehat{Q})=\max_{1\leq k\leq n}\,|a_{(k)}-b_{(k)}|.$
Proof.
[…]
Lemma P.2 (Uniform empirical $W_{\infty}$ reduces to a bottleneck matching problem).
Let $\widehat{P}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\widehat{Q}=\frac{1}{n}\sum_{j=1}^{n}\delta_{y_{j}}$ on $\mathbb{R}^{d}$. For $t\geq 0$, define the bipartite graph $G_{t}$ on $(\{1,\ldots,n\},\{1,\ldots,n\})$ with an edge $(i,j)$ iff $\|x_{i}-y_{j}\|_{2}\leq t$. Then
$W_{\infty}(\widehat{P},\widehat{Q})\leq t\quad\Longleftrightarrow\quad G_{t}\text{ has a perfect matching}.$
Consequently,
$W_{\infty}(\widehat{P},\widehat{Q})=\min\{t\geq 0:\ G_{t}\text{ has a perfect matching}\}.$
Proof.
[…]
Lemma P.3 (Elementary domination for asymptotic sums).
For all $n\geq 3$, one has $n\leq n\log n$ (with $\log$ the natural logarithm). Consequently, for any constants $c_{1},c_{2},c_{3}\geq 0$ and all $n\geq 3$,
$c_{1}dn+c_{2}n\log n+c_{3}n\;\leq\;(c_{1}+c_{2}+c_{3})\,(dn+n\log n).$
Proof.
For $n\geq 3$, $\log n\geq 1$, hence $n\leq n\log n$. Thus $c_{3}n\leq c_{3}n\log n$ and the inequality follows by grouping terms. ∎
Proposition P.1 (Formal Version of Proposition 4.1).
Fix a Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$.
  1. 1.
    Per-slice computation is exact in 1-D. For a fixed direction $u$, the empirical $W_{\infty}$ between the one-dimensional projected samples $\{\langle x_{k},u\rangle\}_{k=1}^{n}$ and $\{\langle y_{k},u\rangle\}_{k=1}^{n}$ equals $\max_{k}|a_{(k)}-b_{(k)}|$ after sorting (Lemma P.1). Per instance, projection costs $O(dn)$ and sorting costs $O(n\log n)$, hence
    $T_{\mathrm{1D}}(u)=O\!\big(\mathsf{M}\,(dn+n\log n)\big).$
  2. 2.
    Finite slicing over $L$ directions. Repeating (1) for $u=U_{\ell}$ over all $\ell\in[L]$ gives
    $T_{\mathrm{slice}}=O\!\big(\mathsf{M}\,L\,(dn+n\log n)\big).$
  3. 3.
    Unsliced $\infty$-Wasserstein. For each instance, forming all distances $\|x_{i}-y_{j}\|_{2}$ requires $\Theta(dn^{2})$ arithmetic and $\Omega(n^{2})$ memory to materialize the distance table. Moreover, exact evaluation of $W_{\infty}$ can be performed by searching over thresholds $t$ and checking perfect-matching feasibility in $G_{t}$ (Lemma P.2). If $\mathcal{T}:=\{\|x_{i}-y_{j}\|_{2}:\ 1\leq i,j\leq n\}$, then $|\mathcal{T}|\leq n^{2}$ and binary search over $\mathcal{T}$ requires $O(\log n)$ feasibility checks. Using a standard bipartite matching algorithm with $O(n^{3})$ time, the per-instance runtime is
    $T_{\mathrm{unsliced,exact}}(1\ \text{instance})=O(dn^{2}+n^{3}\log n),$
    and over all $M$ instances,
    $T_{\mathrm{unsliced,exact}}=O\!\big(\mathsf{M}\,(dn^{2}+n^{3}\log n)\big),$
    with memory $\Omega(n^{2})$ per instance if distances are materialized.
Proof.
We count arithmetic operations and comparisons in the standard word-RAM model. Fix any instance $(\theta,(s_{i},s_{j}))\in\widehat{\Theta}\times\widehat{\mathcal{Q}}$ and let $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ and $y_{1},\ldots,y_{n}\in\mathbb{R}^{d}$ be the corresponding samples.
(1) Per-slice computation with one fixed direction $u$.
Fix $u\in\mathbb{S}^{d-1}$. For each $k\in[n]$, compute $a_{k}:=\langle x_{k},u\rangle$ and $b_{k}:=\langle y_{k},u\rangle$. A dot product in $\mathbb{R}^{d}$ can be computed using exactly $d$ multiplications and d1d-1 additions; thus there exists a constant $c_{\mathrm{dot}}>0$ such that each dot product costs at most $c_{\mathrm{dot}}\,d$ primitive operations. Hence computing all $2n$ projections costs at most $2c_{\mathrm{dot}}\,dn$ operations.
Next, sort $\{a_{k}\}_{k=1}^{n}$ and $\{b_{k}\}_{k=1}^{n}$ to obtain order statistics $a_{(1)}\leq\cdots\leq a_{(n)}$ and $b_{(1)}\leq\cdots\leq b_{(n)}$. Using a standard comparison-based sorting algorithm (e.g., mergesort), sorting $n$ reals costs at most $c_{\mathrm{sort}}\,n\log n$ comparisons for some constant $c_{\mathrm{sort}}>0$ (for all $n\geq 2$). Thus sorting both lists costs at most $2c_{\mathrm{sort}}\,n\log n$ operations.
Define $\widehat{P}_{u}:=\frac{1}{n}\sum_{k=1}^{n}\delta_{a_{k}}$ and $\widehat{Q}_{u}:=\frac{1}{n}\sum_{k=1}^{n}\delta_{b_{k}}$. By Lemma P.1,
$W_{\infty}(\widehat{P}_{u},\widehat{Q}_{u})=\max_{k\in[n]}|a_{(k)}-b_{(k)}|.$
Computing the maximum can be done by one scan through $k=1,\ldots,n$, performing $O(1)$ operations per $k$; hence there exists $c_{\mathrm{scan}}>0$ such that this step costs at most $c_{\mathrm{scan}}\,n$ operations.
Therefore, for this fixed instance and fixed $u$, the total cost is bounded by
$T_{\mathrm{inst}}(u)\;\leq\;2c_{\mathrm{dot}}\,dn\;+\;2c_{\mathrm{sort}}\,n\log n\;+\;c_{\mathrm{scan}}\,n.$
For all $n\geq 3$, Lemma P.3 implies that
$T_{\mathrm{inst}}(u)\;\leq\;C_{\mathrm{inst}}\,(dn+n\log n),\quad C_{\mathrm{inst}}:=2c_{\mathrm{dot}}+2c_{\mathrm{sort}}+c_{\mathrm{scan}}.$
Hence $T_{\mathrm{inst}}(u)=O(dn+n\log n)$.
Now fix $u$ and consider the $\mathsf{M}$ instances. Computing $\Delta_{\infty}^{u}$ requires computing $W_{\infty}(\widehat{P}_{u},\widehat{Q}_{u})$ once per instance and taking their maximum. Maintaining a running maximum costs at most $c_{\max}$ operations per instance for some constant $c_{\max}>0$, thus at most $c_{\max}\mathsf{M}$ operations in total. Consequently, for all $n\geq 3$,
$T_{\mathrm{1D}}(u)\leq\sum_{r=1}^{\mathsf{M}}T_{\mathrm{inst}}^{(r)}(u)\;+\;c_{\max}\mathsf{M}=O\!\big(\mathsf{M}(dn+n\log n)\big),$
which proves (1).
(2) Finite slicing over $L$ directions.
For each $\ell\in[L]$, computing $\Delta_{\infty}^{U_{\ell}}$ is exactly the computation in (1) with $u=U_{\ell}$. Therefore,
$T_{\mathrm{slice}}=\sum_{\ell=1}^{L}T_{\mathrm{1D}}(U_{\ell})=O\!\big(\mathsf{M}L(dn+n\log n)\big),$
which proves (2).
(3) Unsliced $\infty$-Wasserstein (exact computation).
Fix one instance and define $\widehat{P}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\widehat{Q}:=\frac{1}{n}\sum_{j=1}^{n}\delta_{y_{j}}$. Compute the full distance table $D\in\mathbb{R}^{n\times n}$ with entries $D_{ij}:=\|x_{i}-y_{j}\|_{2}$. Computing one Euclidean distance requires $\Theta(d)$ arithmetic operations; hence computing all $n^{2}$ distances requires $\Theta(dn^{2})$ arithmetic operations. If the table is materialized, it stores $n^{2}$ reals and therefore requires $\Omega(n^{2})$ memory.
Let $\mathcal{T}:=\{D_{ij}:i,j\in[n]\}$ be the set of candidate thresholds, with $|\mathcal{T}|\leq n^{2}$. By Lemma P.2,
$W_{\infty}(\widehat{P},\widehat{Q})=\min\{t\in\mathcal{T}:\ G_{t}\ \text{has a perfect matching}\}.$
Sorting the multiset of at most $n^{2}$ values in $\mathcal{T}$ costs $O(n^{2}\log(n^{2}))=O(n^{2}\log n)$, and binary search over the sorted list uses $O(\log|\mathcal{T}|)=O(\log n)$ feasibility checks. Given $t$, scanning $D$ to determine edges of $G_{t}$ costs $O(n^{2})$ time, which is dominated by matching. A perfect matching in a bipartite graph on $2n$ vertices can be found in $O(n^{3})$ time by standard algorithms. Thus each feasibility check costs $O(n^{3})$ time, and the total time spent in feasibility checks is $O(n^{3}\log n)$.
Combining the distance-table computation, threshold sorting, and feasibility checks yields the per-instance bound
$T_{\mathrm{unsliced,exact}}(1\ \mathrm{instance})$
$=O(dn^{2})+O(n^{2}\log n)+O(n^{3}\log n)$
$=O(dn^{2}+n^{3}\log n).$
Over $\mathsf{M}$ instances this becomes $O(\mathsf{M}(dn^{2}+n^{3}\log n))$, with $\Omega(n^{2})$ memory per instance if $D$ is materialized. This proves (3) and completes the proof. ∎
Appendix Q Proof of Proposition 5.1
@@
For any slice $u_{\ell}\in\mathcal{U}$, this implies
$\bigl|\langle\bar{g}_{t}(\mathcal{B}(x;r))-\bar{g}_{t}(\mathcal{B}(x^{\prime};r)),u_{\ell}\rangle\bigr|\leq\frac{2C}{B_{t}}K_{t}(x,x^{\prime};r).$ (64)
@@
By Assumption 5.2, for each $u_{\ell}\in\mathcal{U}$ there exists $L_{t,i}<\infty$ such that, for all $z,z^{\prime}\in\mathbb{R}^{d}$ and all histories $y_{<t}$,
$\bigl|u_{\ell}^{\top}\bigl(T_{t}(z;y_{<t})-T_{t}(z^{\prime};y_{<t})\bigr)\bigr|\leq L_{t,i}\bigl|u_{\ell}^{\top}(z-z^{\prime})\bigr|.$
@@
$\bigl|\langle f_{t}(x,y_{<t};r)-f_{t}(x^{\prime},y_{<t};r),u_{\ell}\rangle\bigr|$ $\leq L_{t,i}\,\frac{2C}{B_{t}}\,K_{t}(x,x^{\prime};r)$
@@
$\bigl|\langle f_{t}(x,y_{<t};r)-f_{t}(x^{\prime},y_{<t};r),u_{\ell}\rangle\bigr|^{2}\leq\Bigl(\frac{2CL_{t,i}}{B_{t}}K_{t}(x,x^{\prime};r)\Bigr)^{2}.$
@@
$\bigl|\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{\ell}\rangle\bigr|^{2}\leq\Bigl(\frac{2K_{t}L_{t,i}C}{B_{t}}\Bigr)^{2}=:h_{t,i}.$
@@
Appendix R Proof of Proposition 5.2
Fix an iteration $t$ and a slice profile $\{\mathcal{U},\omega\}$ with $\mathcal{U}=\{u_{i}\}_{i=1}^{m}\subset\mathbb{S}^{d-1}$. Let $\theta\in\Theta$, $(s_{i},s_{j})\in\mathcal{Q}$, a history $y_{<t}$ in the support, and a coupling $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$ be arbitrary. Let $(X,X^{\prime})\sim\gamma$. In addition, let $r_{t}\sim\mathbb{P}_{\eta,\rho}$ denote the subsampling randomness at iteration $t$. We write $\bar{g}_{t}(X;r_{t})$ and $\bar{g}_{t}(X^{\prime};r_{t})$ for the clipped, averaged minibatch gradients, and $f_{t}(X,y_{<t};r_{t})$, $f_{t}(X^{\prime},y_{<t};r_{t})$ for the corresponding pre-perturbation updates.
Step 1. By per-example $\ell_{2}$-clipping at threshold $C$ and the definition of the discrepancy $K_{t}(X,X^{\prime};r_{t})$, the argument in the proof of Proposition 5.1 gives, for every realization of $r_{t}$,
$\bigl\|\bar{g}_{t}(X;r_{t})-\bar{g}_{t}(X^{\prime};r_{t})\bigr\|_{2}\leq\frac{2C}{B_{t}}K_{t}(X,X^{\prime};r_{t}).$
Hence, for any slice $u_{\ell}\in\mathcal{U}$,
$\bigl|\langle\bar{g}_{t}(X;r_{t})-\bar{g}_{t}(X^{\prime};r_{t}),u_{\ell}\rangle\bigr|\leq\frac{2C}{B_{t}}K_{t}(X,X^{\prime};r_{t}).$ (65)
@@
$\Delta_{t}^{\mathrm{pre}}(X,X^{\prime};r_{t}):=\bar{g}_{t}(X;r_{t})-\bar{g}_{t}(X^{\prime};r_{t}),$
@@
$\Delta_{t}(X,X^{\prime};r_{t}):=f_{t}(X,y_{<t};r_{t})-f_{t}(X^{\prime},y_{<t};r_{t}).$
By Assumption 5.2, for each slice $u_{\ell}\in\mathcal{U}$ and every realization of $r_{t}$,
$\bigl|\langle\Delta_{t}(X,X^{\prime};r_{t}),u_{\ell}\rangle\bigr|\leq L_{t,i}\bigl|\langle\Delta_{t}^{\mathrm{pre}}(X,X^{\prime};r_{t}),u_{\ell}\rangle\bigr|.$
@@
$\bigl|\langle\Delta_{t}(X,X^{\prime};r_{t}),u_{\ell}\rangle\bigr|\leq\frac{2L_{t,i}C}{B_{t}}\,K_{t}(X,X^{\prime};r_{t})$ (66)
@@
Step 2. Squaring both sides of (66) and taking expectations with respect to $\eta_{t}\sim\mathbb{P}_{\eta}$, we obtain, for each $i=1,\dots,m$,
$\mathbb{E}_{\eta_{t}}\bigl[\bigl|\langle\Delta_{t}(X,X^{\prime};r_{t}),u_{i}\rangle\bigr|^{2}\bigr]\leq\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\mathbb{E}_{\eta_{t}}\bigl[K_{t}(X,X^{\prime};r_{t})^{2}\bigr].$
@@
$\mathbb{E}_{\eta_{t}}\bigl[K_{t}(X,X^{\prime};r_{t})^{2}\bigr]\leq\overline{K}_{t}^{2}.$
@@
$\mathbb{E}_{\eta_{t}}\bigl[\bigl|\langle\Delta_{t}(X,X^{\prime};r_{t}),u_{\ell}\rangle\bigr|^{2}\bigr]\leq\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2}=h_{t,i}^{\mathsf{sa}},\quad i=1,\dots,m.$
@@
$h_{t}^{\mathsf{sa}}=\bigl(h_{t,i}^{\mathsf{sa}}\bigr)_{i=1}^{m}\quad\textup{and}\quad h_{t,i}^{\mathsf{sa}}=\Bigl(\frac{2L_{t,i}C}{B_{t}}\Bigr)^{2}\overline{K}_{t}^{2},$
is a valid sa-HUC at iteration $t$. ∎
Appendix S Proof of Lemma 5.1